the average of nine numbers is 40. if 1 is added to the first number, 2 is added to the second number, 3 is added to the third number and so on up to the ninth number, what is the new average?
Answer by RAY100(1637)

(E n )/9 =40
.
{E (n(1) +1) + n(2) +2 +......... } / 9 = ?
.
(E n)/9 + { 1+2+3+4+5+6+7+8+9 } /9 = ?
.
(En)/9 + { 45/9} = ?
.
40 +5 = ?
.
45 = ?